# Real-to-Sim via End-to-End Differentiable Simulation and Rendering

### v1

@@ IV-C World Model Identification
Even with our jointly differentiable physical and appearance models, solving Eqn. 1 can still be rather difficult. This is mainly because our initial guess can be rather noisy and discrepant from the observations. As a result, the naïve gradient descent method can take many iterations and is thus prone to falling into local minima. To mitigate this issue, we propose to use two stages, and we further leverage 3D foundation models trained on web-scale data to generate good initial guesses of all the rigid objects in the scene from partial visual observations. Here we first detail our two-stage optimization formulation and practical considerations for generating reasonable initial guesses.
@@ IV-C1 Two-stage Optimization
We notice that while the initial guess might deviate from our observations in terms of geometry, appearance, and physical parameters, the deviation in geometry and appearance can be largely corrected by considering the first static observation alone, i.e. $\langle{t_{1}},e^{t_{1}},f^{t_{1}},o^{t_{1}},c^{t_{1}},m^{t_{1}}\rangle$. Therefore, our first stage considers only the first time instance and optimize $\theta$ using the following loss:
@@
$c_{3}\mathcal{L}_{\text{pcd}}(o^{t_{1}},\theta)+c_{4}\mathcal{L}_{\text{pen}}(\theta)+c_{5}\sum_{1}^{k}\|q^{i}-q^{t_{1}}\|^{2}+$
$c_{6}\mathcal{L}_{\text{reg}}(\theta,\theta_{0})+c_{7}\mathcal{L}_{\text{smooth}}(\theta),$
@@
with $c_{\bullet}$ being the corresponding weight terms. Here $\hat{o}^{t_{1}}$ is the rendered RGB-D images. $\mathcal{L}_{\text{rgb}}$ is a loss on the RGB images, defined as a weighted sum of $l_{1}$ distance and D-SSIM terms: $\mathcal{L}_{\text{rgb}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{SSIM}}$, where we set $\lambda=0.2$. $\mathcal{L}_{\text{depth}}$ is the $l_{1}$ distance on the depth images. The three loss terms $\mathcal{L}_{1,\text{SSIM},\text{depth}}$ are further masked by $m^{t_{1}}$. $\mathcal{L}_{\text{pcd}}$ is a unilateral Chamfer distance between the observed RGB-D point cloud belonging to the object and the SaP point cloud $\mathcal{P}$, which is defined as the average of the minimum distance between each point on the observed point cloud and any point among the SaP points. The fourth and fifth terms encourage the object to achieve static force equilibrium while having no penetrations, by penalizing the object mesh penetrations in the terrain and the change in object positions over $k$ steps, which is the sum of the positional changes over $k$ steps by simulating forward with no robot actions. These terms provide a strong hint for the occluded part of the object. For example, when an object is lying on the table, our RGB-D observation will not cover the back side of the object. However, our model will guide SaP to fill the back-side geometries by encouraging the object to settle on the table. Finally, the last two terms regularize the object shape, where ${L}_{\text{reg}}$ is the $L_{2}$ norm between the SaP points and the initial SaP points and $L_{smooth}$ is the Laplacian smoothing objective on the object mesh.
@@
We observe that the effects of using $\mathcal{L}_{\text{depth}}$ and $\mathcal{L}_{\text{pcd}}$ complement each other. When only $\mathcal{L}_{\text{depth}}$ is used, we cannot handle the case where a depth pixel on the ground-truth image hits a mesh surface whereas the same pixel is not on the predicted image, in which case there is no gradient information for the geometry to expand and cover this pixel. Similarly, the observed point cloud used in $\mathcal{L}_{\text{pcd}}$ is based on a partial point cloud of the object, and SaP has no information about the boundaries of the object. $\mathcal{L}_{\text{depth}}$ informs our model of the background area and prevents the object geometries from occupying the supposed background.
@@
After the first stage, we have tuned our model to match the first observation. When we move on to the second stage, we incorporate all timesteps by applying the robot pushing forces. For this stage, we use the simpler loss terms: $\mathcal{L}=c_{8}\mathcal{L}_{\text{pcd}}+c_{9}\mathcal{L}_{\text{robot}}$, where the first term have the same definitions as in the first stage and $\mathcal{L}_{\text{robot}}$ is the squared distance between the ground-truth and predicted robot positions. During our optimization, the chain of gradient is back-propagated through the following recursive rule:
@@
where the first two terms ${\partial{\mathcal{L}}}/{\partial{\theta}},{\partial{\mathcal{L}}}/{\partial{q^{i}}}$ is the derivatives of the rendering equation, and the remaining terms in the bracket is the derivatives of the simulator.
@@ IV-C2 Geometry Prior
Our method relies on a reasonable initial guess. Imagine the case with an object settling on the edge of a table and the camera does not observe the contact between the two. The initial guess of the occluded part of the object could be very short and cause the object to directly fall down without touching the table. This cannot be recovered by our optimization since there are no gradients for correcting the geometries. To obtain a reasonable initial guess of the geometries and appearance of the rigid object in interest from partial visual observations, we take advantage of large reconstruction models [32] that predict object 3D models from a single RGB image, trained on web-scale data. In particular, we use TripoSR [33] in our experiments with the segmented RGB image of the object as input image. Since the generated mesh is scale- and transform-agnostic, we subsequently apply RANSAC and scale-aware iterative closest point (ICP) algorithms to register the mesh to the partial object point cloud, computed from the RGB-D image at the first time instance.
@@ V Experiments and Results
Simulation Real World
Figure 3: The experiment setups for the simulated (left) and physical (right) experiments. Two objects (Power Drill from YCB dataset and a box) in the PyBullet simulator are used for simulation. For the real-world experiments, three YCB objects (Power Drill, Mustard Bottle, and Domino Sugar Box) are used. A UR5e arm equipped with a pusher and an overhead Realsense D435 RGB-D camera are used.
Novel View Synthesis Train Dynamics Error Test Dynamics Error
Object MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ Unilateral Chamfer (mm) Pos. (mm) Rot. () $\mu$ Unilateral Chamfer (mm) Pos. (mm) Rot. ()
Box 0.00312 0.950 25.1 1.99 12.2 0.749 0.107 3.50 11.6 23.3
Drill 0.00413 0.942 24.0 4.91 15.1 3.85 0.0970 8.21 12.3 12.7
TABLE I: Novel view synthesis and dynamics prediction errors for the simulated experiments
Train Trajectory3 Test Trajectories
Figure 4: The train (left) and 3 different test (right) trajectories of the Drill object during stage 2 in simulation, with the first and last frames shown. The blue sphere represents a floating robot. The train and testing trajectories are the same for the other objects in both simulated and physical experiments. [Best viewed in color.]
Train TrajectoryTest Trajectory 1Test Trajectory 2Test Trajectory 3
Figure 5: The predicted and ground-truth poses of the Drill object at the last frame of the train and the 3 different test trajectories in one optimization trial. The predicted object poses with the optimized $\theta$ highlighted with a yellow silhouette are overlaid with the ground-truth object, robot, and background. [Best viewed in color.]
View 1View 2 View 3
Figure 6: The ground-truth (left) and predicted (right) RGB images of 3 novel views of the Drill in simulation. Because the optimized mesh shape is slightly larger than the ground truth, the optimized color (with an overhead view during training) at some locations on the mesh has the color of the terrain. The terrain checkers are not as sharp as the ground truth due to the use of point rendering of the colored terrain point cloud.
@@
Figure 7: The initial guess and optimized shape of the Box object in simulated experiments and the Sugar object in physical experiments during stage 1. The initial guess of the Box with the bottom bulges out would penetrate the terrain underneath. On the other hand, the initial mesh of Sugar is too thin and does not touch the terrain. Our algorithm is able to optimize the shape so that it satisfies the physics, albeit not perfect since the objects are completely occluded at the bottom.
V-A Simulated Experiments
@@
Shown in Fig. 3, we conduct all the simulated experiments using data collected with the PyBullet simulator [34]. We use a simple green box object (Box) and a power drill object (drill) from the YCB object dataset [35]. The objects are placed on a flat surface with checker patterns and pushed by a floating sphere robot, while a static overhead camera takes pictures. The optimized simulator parameters are then tested on three different pushes. The pushing trajectories are visualized in Fig. 4 where they are the same for both objects. All the trajectories have $T=30$ time steps with $\delta t=0.01$ s. We use the following weights for optimization: $[c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{9}]=[10,5000,500,1,1000,100,4000,500,100]$. These terms are not carefully tuned and are set such that each term has a similar order of magnitudes at the start of optimization for our experiments. In addition, we set $k=3$ steps for simulating forward with no actions in Stage 1.
The physical parameter we optimize for is the coefficient of friction between the object and the terrain. We assume the mass is known here since optimizing both the mass and the coefficient of friction with a pushing trajectory is ill-conditioned. However, the mass can be easily identified with some prehensile trajectory like picking up the object. We also assume that the center of mass is at the geometry center of the initial geometry guess and that the inertia matrix is calculated assuming the object is a box with the dimensions of the object bounding box. We note that while we make these simplifying assumptions, this is not a limitation of our method and these physical parameters can all be optimized for in the fully differentiable pipeline with more online data.
We report the quantitative results for both objects in Table I, where we show the quality of novel view synthesis and the errors in predicting the dynamics. For each object, we run three optimization trials with different random initial coefficients of friction between 0.1 and 1, and the average across the trials are reported in Table I. We also show the final location of the pushed objects in both the training and testing pushing trajectories in Fig. 5 for one trial and some examples of novel view synthesis in Fig. 6. Our method identifies the shape and appearance parameters, and coefficient of friction that lead to accurate dynamics prediction and high-quality novel view synthesis. We think the error in the coefficient of friction could be due to the inconsistencies between the physics engine of our method and PyBullet. Due to the slight mismatching between the optimized and ground-truth shape of the object geometry, parts of the sides of the optimized Drill mesh have the color of the terrain. We believe this could be resolved with slightly more diverse angles other than a static overhead camera. For the Box object, instead of using an initial geometry guess generated by a reconstruction model, we use an arbitrarily created box that clearly penetrates the terrain underneath to demonstrate the ability of our method to adjust occluded geometry based on the physics, shown in Fig. 7 on the left.
Object Train Unilateral Chamfer (mm) Test Unilateral Chamfer (mm)
Drill 5.22 7.10
Mustard 3.56 6.16
Sugar 4.16 5.26
TABLE II: Dynamics prediction errors for the real world experiments
Train PushTrain TrajectoryTest Trajectory 1Test Trajectory 2Test Trajectory 3
Figure 8: Train and test results of the physical experiments. Each row shows the results from an object. The columns from left to right are: 1) the robot’s push to collect the training data, 2) the optimized object position at the end of the train trajectory, and 3)-5) those at the end of the three different test trajectories. The predicted object poses with the optimized $\theta$ highlighted with a yellow silhouette are overlaid with the ground-truth object and robot, and the background rendered from the optimized simulator. [Best viewed in color.]
@@ V-B Physical Experiments
Shown in Fig. 3, we conduct physical experiments with a UR5e robot arm equipped with an embedded end-effector F/T sensor and a pusher, and a static overhead RealSense D435 RGB-D camera. We adopt the same train and test pushing trajectories as the simulated experiments for 3 YCB objects: power drill (Drill), sugar box (Sugar), and mustard bottle (Mustard). However, we use a total of $T=48$ time steps with $\delta t=0.03$ s. We use the same optimization settings as the simulated experiments.
The dynamics prediction results are reported in Table II, and the train and test prediction results are visualized in Fig. 8. For each object, we use 0.4 as the initial guess of the coefficient of friction for optimization. Note that because we do not have access to the ground-truth object poses and coefficients of friction, we only report the unilateral chamfer distances. Overall, the errors are comparable to those from the simulated experiments. In addition, we show the initial and optimized shape of the Sugar object in Fig. 7 on the right, where the initial shape is too thin and does not contact the terrain below. Our algorithm modifies the occluded geometry to satisfy the physics. While the occluded geometry is not perfect, our method allows the object geometry to be further improved if given more diverse online data.

### v3

@@ IV-C World Model Identification
Even with our jointly differentiable physical and appearance models, solving Eqn. 1 can still be rather challenging. This is mainly because our initial guess can be very poor, especially in the occluded region. As a result, the naïve gradient descent method can take many iterations and is prone to converging to poor local minima. To mitigate this, we use two stages of optimization, and we further leverage 3D foundation models trained on web-scale data to generate good initial guesses of the rigid object in the scene from partial visual observations.
Figure 4: The pushing trajectories used in the experiments. Left: The 8 starting locations of the floating spherical robot pushing trajectories and 3 pushing directions towards the robot at one of the starting locations for the Drill object in the simulation experiments. Middle and right: the training trajectory and 2 sample testing trajectories, with the first and last frames shown. [Best viewed in color.]
Figure 5: The predicted and ground-truth poses of the 5 different objects at the end of sampled testing trajectories for the simulation experiments. After training, the predicted poses are obtained by applying the control forces from the initial pose and integrating forward in time. The predicted object poses are highlighted with a yellow silhouette and overlaid with the ground-truth object, blue floating spherical robot, and background. [Best viewed in color.]
View 1View 2 View 3
Figure 6: The ground-truth (left) and predicted (right) RGB images of 3 novel views of the Drill in simulation. The optimized mesh shape and geometry match the ground truth well, although lacking the fine details that can not be observed from the top view. The terrain checkers are not as sharp as the ground truth due to the use of point rendering of the colored terrain point cloud.
@@ IV-C1 Two-stage Optimization
We note that while the initial guess can deviate significantly from our observations, deviations in geometry and appearance can be largely corrected by considering only the first observation, i.e. $\langle{t_{1}},e^{t_{1}},f^{t_{1}},o^{t_{1}},c^{t_{1}},m^{t_{1}}\rangle$. Therefore, our first stage considers only the first time instance and optimizes $\theta$ using the following loss:
@@
$c_{3}\mathcal{L}_{\text{pcd}}(o^{t_{1}},\theta)+c_{4}\mathcal{L}_{\text{pen}}(\theta)+c_{5}\mathcal{L}_{\text{balance}}(\theta)+$
$c_{6}\mathcal{L}_{\text{reg}}(\theta,\theta_{0})+c_{7}\mathcal{L}_{\text{smooth}}(\mathcal{P}),$
@@
with $(c_{1},\dots,c_{7})$ denoting weight terms. Here, $\hat{o}^{t_{1}}$ is the rendered RGB-D image at $t_{1}$. $\mathcal{L}_{\text{rgb}}$ is a loss on the RGB images, defined as a weighted sum of $l_{1}$ distance and D-SSIM terms: $\mathcal{L}_{\text{rgb}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{SSIM}}$, where we set $\lambda=0.2$. $\mathcal{L}_{\text{depth}}$ is the $l_{1}$ distance on the depth images. When calculating the three loss terms $\mathcal{L}_{1,\text{SSIM},\text{depth}}$, the robot is masked out according to the segmentation mask $m^{t_{1}}$. $\mathcal{L}_{\text{pcd}}$ is a unilateral Chamfer distance between the point cloud generated from the observed RGB-D image pixels belonging to the object and the SaP point cloud $\mathcal{P}$, which is defined as the average of the minimum distance between each point on the observed point cloud and any point among the SaP points. The terms $\mathcal{L}_{\text{pen}}$ and $\mathcal{L}_{\text{balance}}$ encourage the object to achieve static force equilibrium while having no penetrations. $\mathcal{L}_{\text{pen}}$ penalizes the object mesh penetrations in the terrain, and $\mathcal{L}_{\text{balance}}=\sum_{1}^{k}\|p_{o}^{i+t_{1}}-p_{o}^{t_{1}}\|^{2}$ is the sum of the positional changes of the object from the initial position $p_{o}^{t_{1}}$ over $k$ steps by simulating forward with no robot actions. In the case where the initial guess of the object geometry does not come in contact with the terrain at $t_{1}$, $\mathcal{L}_{\text{balance}}$ allows the computation of the gradient information to expand the geometry towards the terrain once the object falls due to gravity and contacts the terrain during the $k$ steps. We use $k=3$ for all the experiments in this paper. These two terms provide a strong hint for the occluded part of the object. For example, when an object is lying on the table, our RGB-D observation will not cover the bottom of the object. However, our model will guide SaP to fill the bottom-side geometries by encouraging the object to settle on the table. Finally, the last two terms regularize the object shape, where ${L}_{\text{reg}}$ is the $L_{2}$ norm between the SaP points and the initial SaP points and $L_{smooth}$ is the Laplacian smoothing objective on the object mesh.
@@
In addition, the use of both $\mathcal{L}_{\text{depth}}$ and $\mathcal{L}_{\text{pcd}}$ is necesary. When only $\mathcal{L}_{\text{depth}}$ is used, if the estimated mesh is smaller than the ground-truth object geometry, the predicted depth pixels that are supposed to reach the ground-truth mesh do not hit the estimated mesh, and there is no gradient information for expanding the geometry. Similarly, the observed point cloud used in $\mathcal{L}_{\text{pcd}}$ is based on a partial point cloud of the object, and SaP has no information about the boundaries of the object. $\mathcal{L}_{\text{depth}}$ informs our model of the background area and prevents the object geometries from occupying the supposed background.
@@
After the first stage, we have tuned our model to match the first observation. When we move on to the second stage, we incorporate all timesteps by applying the robot controls. For this stage, we assume the current estimated geometry is close to the ground truth and use the simpler loss: $\mathcal{L}=c_{8}\mathcal{L}_{\text{pcd}}+c_{9}\mathcal{L}_{\text{robot}}$, where the first term has the same definition as in the first stage and $\mathcal{L}_{\text{robot}}$ is the squared distance between the ground-truth and predicted robot joint angles. Instead of using the observations of every frame for calculating the loss that would be computationally very expensive, we apply the robot control forces from the initial state, integrate forward in time, and render two intermediate and the last frames for calculating the losses. During our optimization, the chain of gradient is back-propagated through the following recursive rule:
@@
where the first two terms ${\partial{\mathcal{L}}}/{\partial{\theta}},{\partial{\mathcal{L}}}/{\partial{q^{i}}}$ is the derivatives of the rendering equation, and the remaining terms in the bracket are the derivatives of the simulator.
Dynamics Parameter Error Trajectory Prediction Error
Method mass (kg) $\mu$ Unilateral Chamfer (mm) Pos. (mm) Rot. () Trans. Vel. (m/s2)
Ours 0.0728 0.106 8.69 15.5 16.7 0.0351
PhysGS [25] 0.225 0.400 24.2 42.8 31.8 0.436
TABLE I: Average Dynamics Parameter Identification and novel trajectory prediction errors for all objects in the simulation experiments
@@ IV-C2 Geometry Prior
Our method relies on a reasonable initial guess. Imagine the case with an object settling on the edge of a table and the camera does not observe the contact between the two. The initial guess of the occluded part of the object could be very short and cause the object to directly fall down without touching the table. This cannot be recovered by our optimization since the object never hits the table and there are no gradients for correcting the geometries. To obtain a reasonable initial guess of the geometries and appearance of the rigid object of interest from partial visual observations, we take advantage of large reconstruction models [34] that predict object 3D models from a single RGB image, trained on web-scale data. In particular, we use TripoSR [35] in our experiments with the segmented RGB image of the object as the input image. Since the generated mesh is scale- and transform-agnostic, we subsequently apply RANSAC and scale-aware iterative closest point (ICP) algorithms with Open3D [36] to register the mesh to the partial object point cloud, computed from the RGB-D image at the first time instance.
@@ V Experiments and Results
Figure 7: Results of three example testing trajectory of the physical experiments. The predicted poses are obtained by applying the control forces from the initial pose and integrating forward in time. The predicted object poses with the optimized $\theta$ highlighted with a yellow silhouette are overlaid with the ground-truth object and robot, and the background rendered from the optimized simulator. [Best viewed in color.]
@@
Figure 8: The initial guess and optimized shape of the Box object in simulated experiments and the Sugar object in physical experiments during stage 1. The algorithm is able to correct the bulging on the underside of the Box that would intersect the terrain. On the other hand, the initial mesh of Sugar is too thin and does not touch the terrain. Our algorithm is able to optimize the shape so that it satisfies physics constraints.
V-A Simulation Experiments
@@
Shown in Fig. 3, we conduct all the simulated experiments using data collected with the PyBullet simulator [37]. We use a simple green box object (Box) and 8 objects (Gelatin, RubiksCube, Spam, TunaCan, Mustard, Bleach, Drill, and Sugar) from the YCB object dataset [38]. The objects are placed on a flat surface with checker patterns and pushed by a floating sphere robot, while a static overhead camera takes pictures. The world model parameters are first optimized with data collected on one training push, and evaluated on 23 novel pushes that are drastically different. Fig. 4 shows the starting positions and the pushing directions of the pushes for the Drill object. Similar pushes are used for all the other objects where the starting locations are adjusted based on the size of the objects. The trajectories push the objects up to 12 cm and 80. All the trajectories have $T=30$ time steps with $\delta t=0.01$ s. We use the following weights for optimization: $[c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{9}]=[10,5000,500,1,1000,100,4000,500,100]$. These terms are not carefully tuned and are set such that each term has a similar order of magnitude at the start of optimization for our experiments. The physical parameters we optimize for include the mass, surface coefficient of friction, the center of mass, and the rotational inertial. The center of mass is initialized at the geometry center of the initial shape guess. The rotational inertia is initialized by treating the object as a box, whose dimensions are the bounding box of the initial geometry guess. We assume that the rotational inertia only has diagonal terms. For all experiments including simulations and real-world experiments, the surface coefficient of friction is initialized at 0.2 and the mass is initialized at 0.2 kg.
We first evaluate whether our method can identify the shape and physical parameters accurately such that it generalizes well to new physical interactions. We compare our method against a recently proposed method that represents the world jointly with Gaussian Splats and physical particles and allows it to perform both novel-view rendering and physics-based trajectory predictions. We refer to this method as PhysGS. This method performs simulation with a particle-based simulator [39]. In the original paper, the physical parameters of the particles are arbitrarily set, and virtual forces are optimized to match the observations and predictions on the training trajectory. To allow more accurate prediction for new trajectories, we optimize the total mass and coefficient of friction of the object particles with grid search using the partial Chamfer distance between the observed object point cloud and the predicted object particles on the last time frame in the trajectory.
We report the quantitative results for dynamics parameter estimation and novel trajectory predictions for all objects in the simulation experiments in Table I, which are the average across all objects. We also shown some qualitative examples that are representative of the average errors in Fig. 5. The dynamics parameter estimation error from the training trajectory, and the average pose error, unilateral chamfer distance, and the translational velocity error at the end of the testing trajectories are reported. The final pose and velocity of the objects are obtained by applying the control forces and integrating forward in time. Our method identifies the dynamics parameters accurately and shows low trajectory prediction errors. We would like to point out in particular that the average rotation error is heavily skewed by the TunaCan object since our method currently uses the object surface point cloud to track the object during stage 2 optimization, and cannot properly differentiate the rotation of a cylindrical object. As a result, the average rotation error is 28.0 for TunaCan, and we aim to address this issue in future work. On the other hand, PhysGS generalizes very poorly to the testing trajectories. While this is partially because of the lack of proper dynamics parameters and physics-based shape estimation, we also find that the particle-based simulator is extremely sensitive to simulator parameters and have poor physical fidelity, especially for rigid objects. We also show an example of the initial and optimized geometries of the Box object in Fig. 8 (left), which demonstrate the ability of our method to adjust occluded geometry based on the physics.
Next, we also evaluate the quality of the novel-view synthesis of our method. For each of the testing objects, we evaluate the synthesized RGB images from 10 novel viewpoints around the scene, and our method achieves 0.00225 of mean squared error (MSE), 0.965 of structural similarity index measure (SSIM) and 26.5 of peak signal-to-noise ratio (PSNR). We also show some examples of novel view synthesis of the Drill object in Fig. 6, which matches the ground truth very well.
@@ V-B Physical Experiments
Shown in Fig. 3, we conduct physical experiments with a UR5e robot arm equipped with an embedded end-effector F/T sensor and a pusher, and a static overhead RealSense D435 RGB-D camera. We use one training trajectory and three testing pushing trajectories that all share the same starting location but are spaced about 15 apart as the simulated experiments for 3 YCB objects: power drill (Drill), sugar box (Sugar), and mustard bottle (Mustard). The trajectories push the object up to 9 cm and 30. We use a total of $T=48$ time steps with $\delta t=0.03$ s. We use the same optimization settings as the simulated experiments.
On average across all testing trajectories and testing objects, our method achieves a mass identification error of 0.257 kg and 5.66 mm of unilateral chamfer distance between the observed object point cloud and the predicted object at the last frame of the trajectory. Note that we only reported the mass identification error because the ground truth is very easy to measure while measuring the surface friction requires a specialized setup. Since we do not have access to the ground-truth object poses and only have access to the raw observations, we report the unilateral chamfer distances between the observed object point cloud and the predicted object geometry. The train and test prediction results for three sample trajectories are visualized in Fig. 7. Overall, the errors are comparable to those from the simulated experiments. In addition, we show the initial and optimized shape of the Sugar object in Fig. 8 on the right, where the initial shape is too thin and does not contact the terrain below. Our algorithm modifies the occluded geometry to satisfy the physics.